if total autonomous spending is 800, the marginal propensity to save is 0.2 and the marginal tax rate is 0.25. what is the equilibrium income
1
Expert's answer
2014-05-28T13:15:31-0400
We can write the following formula:

Y= C = Ao + c’(Y-T),
where
Y– equilibrium income
(Y-T)– disposable income
c’– propensity to consume
C– consumption

So,
Y= 800 + (1-0.2)*(Y – 0.25Y) = 800 + 0.8*0.75Y = 800 + 0.6Y
0.4Y= 800
Y= 2000